1.Some students were asked if they carry a credit card. Here are the responses.
Class Credit Card Carrier Not a Credit Card Carrier Total
Freshman 33 27 60
Sophomore 6 34 40
Total 39 61 100
What is the probability that the student is a sophomore and doesn't carry a credit card?
2.Some students were asked if they carry a credit card. Here are the responses.
Class Credit Card Carrier Not a Credit Card Carrier Total
Freshman 33 27 60
Sophomore 6 34 40
Total 39 61 100
What is the probability that the student is a sophomore given he doesn't carry a credit card?

Answer by edjones(7569)

1. 34/100 = 17/50 = .34
.
2. 34/61 = .557
.
Ed